OpenStudy (anonymous):

if AB=2x^2+3x-5,BC=x^2+5x-4,and AC=x^2,solve for x and find AC

hartnn (hartnn):

is B the mid-point of A and C ?

OpenStudy (anonymous):

No 

hartnn (hartnn):

ok, so u have AB=BC so, 2x^2+3x-5 = x^2+5x -4

hartnn (hartnn):

subtraxt x^2 from both sides, what u get ?

OpenStudy (anonymous):

1x^2+3x-5=5x-4

hartnn (hartnn):

subtract 5x and add 4 on both sides

OpenStudy (anonymous):

1x^2-2X-1=0

hartnn (hartnn):

now can u solve this quadratic ?

OpenStudy (anonymous):

Yes thx'

OpenStudy (anonymous):

Sorry No

hartnn (hartnn):

is the question correct ? verify it

OpenStudy (anonymous):

i got x=1+√2 x=1-√2

hartnn (hartnn):

b^2-4ac is incorrect....

OpenStudy (anonymous):

so what are you tring to say did i do it wrong...

hartnn (hartnn):

x^2-2x-1 = 0 did u use quadratic formula ?

OpenStudy (anonymous):

\[x=-b+-\sqrt{b^2-4ac}\]

OpenStudy (anonymous):

that formula?

hartnn (hartnn):

so what is b^2-4ac there ?

hartnn (hartnn):

oh wait,

hartnn (hartnn):

x=1+√2 x=1-√2 is correct

OpenStudy (anonymous):

how do i plug that in to find out what AC is

hartnn (hartnn):

(1+√2 )^2 = 1+2+2√2 = 3+2√2

hartnn (hartnn):

or (1-√2 )^2 = 1+2-2√2 = 3-2√2

hartnn (hartnn):

ok?

OpenStudy (anonymous):

ok